An iron weighs 2.225g. It’s dissolved in an acid which oxidizes to Fe2+ and made up of 250cm3. A 25cm3 of sample required is 31cm3 of a 0.0185mol L-1 solution of potassium dichromate in a titration.

Calculate the number of moles of potassium dichromate used in the titration?


1
Expert's answer
2021-08-09T07:47:36-0400

Molar Mass of iron = 55.485

Molar Mass of potassium dichromate =294.185

= 18.5 × 31 = 573.5

= 573.5/294.185

= 1.95moles